OpenStudy (anonymous):

Suppose you wish to have $15,000 in 19 years. Use the present value formula to find how much you should invest now at 9% interest, compounded annually in order to have $ 15,000, 19 years from now. a. $2,917.35 b. $1,350.00 c. $12,082.65 d. $9,165.30

OpenStudy (whpalmer4):

present value formula is \[F = P(1+i)^n\]where \(F\) is future value, \(P\) is present value, \(i\) is interest rate expressed as a decimal, \(n\) is the number of compounding periods We have \(F=15000\), \(n=19\), \(i = 0.09\). Plug in the numbers and solve for \(P\)

OpenStudy (anonymous):

How do I do that

OpenStudy (whpalmer4):

Well, \[F=P(1+i)^n\]We want to have P all alone with everything else on the other side of the = sign. Let's divide both sides by \((1+i)^n\) \[\frac{F}{(1+i)^n} = P\]Okay, now we just plug in the numbers: \[P = \frac{15000}{(1+0.09)^{19}} = \frac{15000}{1.09^{19}}=\]

OpenStudy (anonymous):

so I divide them

OpenStudy (whpalmer4):

first you have to find the value of 1.09^19. then divide 15000 by that value.

OpenStudy (anonymous):

5.141661255

OpenStudy (whpalmer4):

good. what is 15000/5.141661255?

OpenStudy (anonymous):

2917.345048